Project

General

Profile

Actions

Ops/Comm Scheduling meeting

Location : CR11 R3 @ 15:30 pm on Tuesday 1st

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda

  • Actions from last meeting
    • Craig: SSU oxygen monitor needs connecting
    • Melissa/Ed: Define run settings with helium in one tracker and air in the other. Pass these on to Steve B., Chris R. and MOM. Update DC on tracker requirements for this run.
    • Pierrick: Investigate the 'start/stop' nature of the neutron monitor IOC
    • Craig: Liaise with Mark Arnold about scheduling BPS tests in shutdown
    • John G: Arrange for TOF1, TOF2/KL/EMR surveying
    • Paolo: Fix missing archiver data problems
  • MOM report - Data-taking this week
    • Lessons learnt during the run
    • Summary of runs (MOM)
    • Data plots from recent shift
    • Record keeping in the control room
      • Concerns raised about elog and paper log record keeping
      • Software change control procedure must be followed during data-taking.
  • DC report - activities in the Hall this week
  • Decay Solenoid PS installation and calibration
    • Kaeser compressor status
    • PSU commissioning status
  • Focus coil and absorber installation status
    • Readiness to install LiH absorber
  • Controls and monitoring from last meeting
    • Radiation monitor EPICS display freezes for ~2min
    • New controls and monitoring issues?
  • Scheduling of required hall work and operations
    • Installation of LiH
    • Pre-planning for SSU/FC/SSD powering
  • AOB
  • Chair of meetings for next 3 weeks
    • Steve is teaching at this time for the next 3 weeks. DC will chair the meeting.
  • DONM : 8/3/2016

Minutes

Present: Ryan, Ed, Henry, Melissa, Pierrick, Andy G, Paul K, Paolo, Durga, Francois, Colin, Joe, Vicky B, Victoria B, Ken, Chris Hunt, Josef, John G, Mark T, Craig, Jaroslaw, Chris R, Mike C.

  • Actions from last meeting
    • Craig: SSU oxygen monitor needs connecting
      • Done, but some outstanding issues with it
    • Melissa/Ed: Define run settings with helium in one tracker and air in the other. Pass these on to Steve B., Chris R. and MOM. Update DC on tracker requirements for this run.
      • Stands.
    • Pierrick: Investigate the 'start/stop' nature of the neutrino monitor IOC
      • Pierrick thinks this is a problem with the neutron monitor itself
    • Craig: Liaise with Mark Arnold about scheduling BPS tests in shutdown
      • Stands. Also need to include Dean Adams
    • John G: Arrange for TOF1, TOF2/KL/EMR surveying
      • Done, survey will be carried out Wed/Thurs morning. Check at next meeting, results need uploading as a MICE note.
    • Paolo: Fix missing archiver data problems
      • Done
  • Actions arising from this meeting
    • Detector owners: Present one plot at this meeting that shows whether your detector is currently OK/not OK.
    • John G/Colin/Andy: RAMS for LiH installation, discuss requirements.
    • Steve B/Chris R/MOM: Determine amount of empty and full absorber data required. When is the best time to install the LiH?
    • MOM/DC: Co-ordinate the best time to change over from empty to LiH absorber -- default date is Monday 7th March.
    • Jaroslaw: Determine ramping/training plan for SSU/FC/SSD in consultation with Colin, Ken, Chris R, Pierrick, Alan B, Mark P, Josef
    • DC: Reintroduce the dummy for searches
  • Notes from this meeting
    • DS PSU installation: Got to 120A before it tripped. All fine up until then. Circuit breakers OK. Need to understand the instability at 120A. Expert required to start up the DS from 0A. There will be a soak test.
    • Request from Mike C: please try to co-ordinate system reboots so that they all happen together.
    • DS compressor has a very small leak. Monitor for now, but arrange for service in shutdown.
    • LiH: Concerns about the amount of time it will take to install the absorber without exposing it to water vapour. Discussion about using argon gas to surround it during installation (this would constitute a change to the risk assessment Andy N. is working on)
    • Searching the MICE Hall must be taken more seriously, particularly when searching around the open PRY.
    • Henry notes that the max currents for the quads are listed on the web whiteboard and must be adhered to.

Updated by Blackmore, Victoria over 6 years ago · 15 revisions